OpenStudy (joshoyen):

If one student can complete the decorations for the prom in 5 days working alone, another student could do it in 3 days, and a third could do it in 4 days, how long would it take them working together?

OpenStudy (michele_laino):

If I call with \(D\) the decoration, then the working rates of the three students, are: \[\Large \frac{D}{5},\;\frac{D}{3},\;\frac{D}{4}\] respectively. So, if they work together, then after \(t\) hours, the work \(D\) is completed, and \(t\) is such that: \[\Large t\left( {\frac{D}{5} + \frac{D}{3} + \frac{D}{4}} \right) = D\] please solve that equation for \(t\)

OpenStudy (michele_laino):

oops.. \(t\) is the number of days

OpenStudy (joshoyen):

ohhh okay thanks!